OpenStudy (anonymous):

Find the angle between the straight lines- \[xcos \alpha +ysin \alpha = p\] \[xcos \beta +ysin \beta = q\]

hartnn (hartnn):

I hava method, donno whether its the best... find slope of each line, m1,m2 angle between lines is = tan^-1 ((m1-m2)/(1+m1m2))

OpenStudy (anonymous):

i know that, but how to find the slope.

hartnn (hartnn):

isolate y slope of y=mx+c is m

hartnn (hartnn):

i think u can isolate y

OpenStudy (anonymous):

ok got it

hartnn (hartnn):

good :)

OpenStudy (anonymous):

is it right ? \[\tan^{-1}( (\cot \beta -\cot \alpha) / (1 +\cot \alpha \cot \beta))\]

OpenStudy (anonymous):

yes